Grade 12th passAlgebra

if f(x) satisfies the relation 2f(x)+f(1-x)=x ^2 for all real x, then find f(x)

twisha
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
We have equation
 
2 f(x) + f(1 -x) = x²       ...(i)
 
Now replace x by (1 - x)
 
2 f(1 -x) + f(x) = (1-x)² ....(ii)
 
Now multiply equation (i) by 2 and eliminate f(1-x)
 
Hence now you will get f(x)